OpenStudy (anonymous):

y-(y^2)/4 = u so y = ?

OpenStudy (anonymous):

Did u write it complate ???

OpenStudy (anonymous):

y^2/4 - y + u = 0 solve quadratic equation for y

OpenStudy (anonymous):

Hi @Coolsector is it y = \[2 \pm 2\sqrt{1-u}\]

OpenStudy (anonymous):

Use from delta and 2th equation form :) Do u know that ?:)

OpenStudy (anonymous):

Hi E.ali I graduated from high school 13 years ago, basically I forgot all this kind of stuff

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Hi and no problem! Look if we have this equation and want to find X : ax^2+bx+C=0 Have this : We have two answer there X1and X2 (1,2are just name of answers .) And to find them : -b+_Delta(under Radical)/2a Delta=b^2-4ac Remember that ?